# Entropic curvature not comparable to other curvatures – or is it?

### v1

@@ 4 Universal lower curvature bounds and perturbation
In this section, we derive explicit lower curvature bounds for general finite Markov chains and present perturbation results. It was shown in [23, Theorem 4.1] that the entropic curvature of any finite Markov chain is always greater than $\infty$- ∞. Our explicit lower curvature bounds are given in Theorem 4.2 for infinite dimension and Theorem 4.3 for finite dimension. While the second result is based on an inequality for the logarithmic mean (Lemma 2.3) and holds therefore only for entropic curvature, the first one is a lower curvature bound for arbitrary means. These explicit results inspired also the derivation of a lower bound for the intertwining curvature given in [28]. The perturbation results are concerned with comparisons of two Markov chains with closeby parameters and closeby means: Theorem 4.7 compares so-called $\delta$-curvatures of two Markov chains with the same graph structure and bounds on the $\log$-differences of their steady states, transition probabilities and means. Theorem 4.8 assumes that both Markov chains are paired with the same mean $\theta$ and provides a bound on the difference between the $\theta$-curvatures of two weighted graphs with the same topology and bounds on the $\log$-differences of the vertex measures and edge weights. All results are given in terms of the minimum $Q_{\min}$ of the positive transition probabilities of the involved Markov chains.
@@ 4.2 Lower entropic curvature bound for finite dimension
$\left(Q_{\min}b_{0}(\rho_{z},\rho_{y},\rho_{x})+(2K-1)\widehat{\rho}_{xy}\right)\left(Q_{\min}b_{0}(\rho_{x},\rho_{y},\rho_{z})+(2K-1)\widehat{\rho}_{zy}\right)\\ \geq\left(-2(\widehat{\rho}_{xy}+\widehat{\rho}_{zy})+\frac{4\rho_{y}}{N}\right)^{2}.$
@@
$(2K-1)Q_{\min}(b_{0}(\rho_{z},\rho_{y},\rho_{x})\widehat{\rho}_{zy}+b_{0}(\rho_{x},\rho_{y},\rho_{z})\widehat{\rho}_{xy})\\ \geq\left(-2(\widehat{\rho}_{xy}+\widehat{\rho}_{zy})+\frac{4\rho_{y}}{N}\right)^{2}.$
@@
$(2K-1)Q_{\min}(b_{0}(\rho_{z},\rho_{y},\rho_{x})\widehat{\rho}_{zy}+b_{0}(\rho_{x},\rho_{y},\rho_{z})\widehat{\rho}_{xy})\\ \geq 8((\widehat{\rho}_{xy})^{2}+(\widehat{\rho}_{zy})^{2})+\frac{16\rho_{y}^{2}}{N^{2}}.$
@@
$\left(8+\frac{32}{N^{2}}\right)(b_{0}(\rho_{z},\rho_{y},\rho_{x})\widehat{\rho}_{zy}+b_{0}(\rho_{x},\rho_{y},\rho_{z})\widehat{\rho}_{xy})\\ \geq 8((\widehat{\rho}_{xy})^{2}+(\widehat{\rho}_{zy})^{2})+\frac{16\rho_{y}^{2}}{N^{2}},$
@@ 4.3 Perturbation results
The requirement $|\nabla\log\rho|<\delta$ in the above definition has been proved useful in the discrete setting. It has been called the regularization trick in [38], and has been used there to compare the log-Sobolev constant with its modified counterpart (see Theorem 1 in [38]). Moreover, we have
@@
[…]
@@
[…]
@@
[…]
@@
$\frac{1}{2}\langle P\rho,\Gamma_{\rho}f\rangle_{\pi}\geq\frac{1}{3e^{3\epsilon}-2}\left(\left(\widetilde{K}+\frac{1}{2}\right)-\left|\widetilde{K}+\frac{1}{2}\right|\left(e^{3\epsilon}-1\right)-\frac{16(e^{3\epsilon}-1)}{Q_{\rm{min}}}\right)\mathcal{A}_{\rho}(f)\\ +\mathcal{A}_{\rho}(f,\Delta f).$
@@ 5 Examples
In this section, we present examples of finite Markov chains $(X,Q,\pi)$ with very different behaviours with regards to their entropic curvatures and Bakry-Émery and Ollivier Ricci curvatures. To simplify the presentation, we drop the requirement of $\pi$ and $\rho$ being normalised. Note, however, that we always require $\sum_{y}Q(x,y)=1$ for our transition probabilities. We also present comparisons with Ollivier’s sectional curvature [34] (see also [25]). Let us briefly recall Ollivier curvature notions.
@@
The Ollivier Ricci curvature of $(X,Q,\pi)$ (see [34]) is defined as follows on the edges $x\sim y$ of $G$:
@@
where $W_{1}$ is the $L^{1}$-Wasserstein distance for probability measures on $X$.
The Ollivier sectional curvature of $(X,Q,\pi)$ (see [32, Problem P]) is defined as follows on the edges $x\sim y$ of $G$:
@@ A Markov chain $(X,Q,\pi)$ with $X=\{0,1,\ldots,n-1\}$ is a birth-death process if
We recall the formulae for computing Ollivier Ricci curvature and Bakry-Émery curvature of birth-death processes, which were given in [29, Theorem 2.10] and [14, Section 2.1], respectively. We remark further that these formulae are also applicable for weighted cycle graphs of length at least $6$.
Theorem 5.3 ([29, Theorem 2.10]).
@@ 5.1 Weighted $3$-point graph
In the following $\alpha$-parametrized weighted three-point birth-death chain $X$ with $X=\{x,y,z\}$ where $x\sim y$ and $y\sim z$, the entropic curvature tends to $\infty$- ∞ (as the parameter $\alpha\to\infty$), while both Bakry-Émery curvature and Ollivier Ricci curvature are strictly positive and bounded away from zero uniformly. Note, however, the Ollivier sectional curvature in this example is $-1$. This example was already considered as a counterexample of the Peres-Tetali conjecture about the log-Sobolov inequality for positive Ollivier curvature (see [26]).
@@ 5.3 Layer-rescaled prism $C_{n}\times K_{2}$
Next, let us recall the Ollivier Ricci curvature formula from [29, Theorem 2.1]:
$K_{ORC}(x,y)=\inf_{\begin{subarray}{c}f\in Lip(1)\\ f(y)-f(x)=1\end{subarray}}(\Delta f(x)-\Delta f(y))$ (52)

### v3

@@ 4 Universal lower curvature bounds and perturbation
In this section, we derive explicit lower curvature bounds for general finite Markov chains and present perturbation results. It was shown in [Mi-13, Theorem 4.1] that the entropic curvature of any finite Markov chain is always greater than $\infty$- ∞. Our explicit lower curvature bounds are given in Theorem 4.2 for infinite dimension and Theorem 4.3 for finite dimension. While the second result is based on an inequality for the logarithmic mean (Lemma 2.3) and holds therefore only for entropic curvature, the first one is a lower curvature bound for arbitrary means. These explicit results inspired also the derivation of a lower bound for the intertwining curvature given in [MWZ-24]. The perturbation results are concerned with comparisons of two Markov chains with closeby parameters and closeby means: Theorem 4.7 compares so-called $\delta$-curvatures of two Markov chains with the same graph structure and bounds on the $\log$-differences of their steady states, transition probabilities and means. Theorem 4.8 assumes that both Markov chains are paired with the same mean $\theta$ and provides a bound on the difference between the $\theta$-curvatures of two weighted graphs with the same topology and bounds on the $\log$-differences of the vertex measures and edge weights. All results are given in terms of the minimum $Q_{\min}$ of the positive transition probabilities of the involved Markov chains.
@@ 4.2 Lower entropic curvature bound for finite dimension
$\left(Q_{\min}b_{0}(\rho_{z},\rho_{y},\rho_{x})+(2K-1)\widehat{\rho}_{xy}\right)\left(Q_{\min}b_{0}(\rho_{x},\rho_{y},\rho_{z})+(2K-1)\widehat{\rho}_{zy}\right)\geq\left(-2(\widehat{\rho}_{xy}+\widehat{\rho}_{zy})+\frac{4\rho_{y}}{N}\right)^{2}.$
@@
$(2K-1)Q_{\min}(b_{0}(\rho_{z},\rho_{y},\rho_{x})\widehat{\rho}_{zy}+b_{0}(\rho_{x},\rho_{y},\rho_{z})\widehat{\rho}_{xy})\geq\left(-2(\widehat{\rho}_{xy}+\widehat{\rho}_{zy})+\frac{4\rho_{y}}{N}\right)^{2}.$
@@
$(2K-1)Q_{\min}(b_{0}(\rho_{z},\rho_{y},\rho_{x})\widehat{\rho}_{zy}+b_{0}(\rho_{x},\rho_{y},\rho_{z})\widehat{\rho}_{xy})\geq 8((\widehat{\rho}_{xy})^{2}+(\widehat{\rho}_{zy})^{2})+\frac{16\rho_{y}^{2}}{N^{2}}.$
@@
$\left(8+\frac{32}{N^{2}}\right)(b_{0}(\rho_{z},\rho_{y},\rho_{x})\widehat{\rho}_{zy}+b_{0}(\rho_{x},\rho_{y},\rho_{z})\widehat{\rho}_{xy})\geq 8((\widehat{\rho}_{xy})^{2}+(\widehat{\rho}_{zy})^{2})+\frac{16\rho_{y}^{2}}{N^{2}},$
@@ 4.3 Perturbation results
The requirement $|\nabla\log\rho|<\delta$ in the above definition has been proved useful in the discrete setting. It has been called the regularization trick in [STY-23], and has been used there to compare the log-Sobolev constant with its modified counterpart (see Theorem 1 in [STY-23]). Moreover, we have
@@
$\left\|\log\frac{\widetilde{Q}}{Q}\right\|_{\infty}\leq\epsilon\quad\text{and}\quad\left\|\log\frac{\widetilde{\pi}}{\pi}\right\|_{\infty}\leq\epsilon,$
@@
$\left|\log\frac{\partial_{1}\widetilde{\theta}(s,1)}{\partial_{1}\theta(s,1)}\right|\leq\epsilon\quad\text{whenever}\quad s\in\left[e^{-\delta},e^{\delta}\right].$ (44)
@@
$\sup_{x\sim y}\left|\log\frac{\widetilde{w}_{xy}}{w_{xy}}\right|\leq\epsilon\quad\text{and}\quad\sup_{x\in X}\left|\log\frac{\widetilde{m}(x)}{m(x)}\right|\leq\epsilon.$
@@
$\frac{1}{2}\langle P\rho,\Gamma_{\rho}f\rangle_{\pi}\geq\frac{1}{3e^{3\epsilon}-2}\left(\left(\widetilde{K}+\frac{1}{2}\right)-\left|\widetilde{K}+\frac{1}{2}\right|\left(e^{3\epsilon}-1\right)-\frac{16(e^{3\epsilon}-1)}{Q_{\rm{min}}}\right)\mathcal{A}_{\rho}(f)+\mathcal{A}_{\rho}(f,\Delta f).$
@@ 5 Examples
In this section, we present examples of finite Markov chains $(X,Q,\pi)$ with very different behaviours with regards to their entropic curvatures and Bakry-Émery and Ollivier Ricci curvatures. To simplify the presentation, we drop the requirement of $\pi$ and $\rho$ being normalised. Note, however, that we require $\sum_{y}Q(x,y)=1$ for our transition probabilities. We also present comparisons with Ollivier’s sectional curvature [Oll-09] (see also [Mu-23]). Let us briefly recall Ollivier curvature notions.
@@
The Ollivier Ricci curvature of $(X,Q,\pi)$ (see [Oll-09]) is defined as follows on the edges $x\sim y$ of $G$:
@@
where $W_{1}$ is the $L^{1}$-Wasserstein distance for probability measures on $X$, defined as
$W_{1}(\mu,\nu):=\inf_{\gamma:X\times X\to[0,1]}\left\{\sum_{u,v\in V}d(u,v)\gamma(u,v)\,\middle|\,\sum_{v}\gamma(u,v)=\mu(u)\,\text{ and }\,\sum_{u}\gamma(u,v)=\nu(v)\right\}.$
The Ollivier sectional curvature of $(X,Q,\pi)$ (see [Oll-10, Problem P]) is defined as follows on the edges $x\sim y$ of $G$:
@@
Ollivier Ricci curvature $K_{ORC}(x,y)$ can also be computed alternatively via the following formula from [MW-19, Theorem 2.1]:
$K_{ORC}(x,y)=\inf_{\begin{subarray}{c}f\in Lip(1)\\ f(y)-f(x)=1\end{subarray}}(\Delta f(x)-\Delta f(y)).$ (48)
@@ A Markov chain $(X,Q,\pi)$ with $X=\{0,1,\ldots,n-1\}$ is a birth-death process if
We recall the formulae for computing Ollivier Ricci curvature and Bakry-Émery curvature of birth-death processes, which were given in [MW-19, Theorem 2.10] and [HM-23, Section 2.1], respectively. We remark further that these formulae are also applicable for weighted cycle graphs of length at least $6$.
Theorem 5.3 ([MW-19, Theorem 2.10]).
@@ 5.1 Weighted $3$-point graph
In the following $\alpha$-parametrized weighted three-point birth-death chain $X$ with $X=\{x,y,z\}$ where $x\sim y$ and $y\sim z$, the entropic curvature tends to $\infty$- ∞ (as the parameter $\alpha\to\infty$), while both Bakry-Émery curvature and Ollivier Ricci curvature are strictly positive and bounded away from zero uniformly. Note, however, the Ollivier sectional curvature in this example is $-1$. This example was already considered as a counterexample of the Peres-Tetali conjecture about the log-Sobolov inequality for positive Ollivier curvature (see [munch2023ollivier]).
@@ 5.3 Layer-rescaled prism $C_{n}\times K_{2}$
Next, we use formula (48) for the computation of Ollivier Ricci curvature.
@@
5.4 Bernoulli-Laplace models
In this subsection we summarize curvature considerations and the Modified Logarithmic Sobolov Inequality (MLSI) for Bernoulli-Laplace models from the literature. We start with the definition of Bernoulli-Laplace models.
Definition 5.5.
A Bernoulli-Laplace model is a finite Markov chain $(X,Q,\pi)$ with the following properties: $X$ is the set of subsets of $[L]:=\{1,\dots,L\}$ of cardinality $N$ with $11\leq N\leq L-$ 1. The transition rates $Q$ (and the associated unique steady state $\pi$) are defined via a collection $\lambda_{j}\geq 0$, $j\in[L]$ of “intensities”, as follows: For $x,y\in X$, $x\neq y$, we define
$Q(x,y)=\begin{cases}\lambda_{j}/L,&\text{if $x\setminus y=\{j\}$ for some $j\in[L]$,}\\ 0,&\text{otherwise,}\end{cases}$
and, for $x\in X$
$Q(x,x)=1-\sum_{y\neq x}Q(x,y).$
The intensities are assumed to be small enough such that
$\inf_{x\in X}Q(x,x)\geq\frac{1}{2}.$
Stochastically, a Bernoulli-Laplace model represents an interacting particle system with $N$ particles distributed on $L$ sites with the constraint that each site can contain at most one particle and where the particle on site $i$ jumps to an unoccupied sites with jump rate only depending on the $\lambda_{i}$.
Note that the underlying combinatorial graph of a Bernoulli-Laplace model is the Johnson Graph $J(L,N)$. Bakry-Émery curvature for the unweighted Johnson graph was given in [KKRT-16] and [CLP-19].
To state the Modified Logarithmic Sobolev Inequality, we need to introduce the following functionals for $f,g\in\mathbb{R}^{X}$ and $\rho\in\mathcal{P}_{*}(X)$ (that is, $\rho>0$ with $\sum_{x\in X}\rho(x)\pi(x)=1$):
$\mathcal{E}(f,g)$ $:=\langle\nabla f,\nabla g\rangle_{\pi}=\frac{1}{2}\sum_{x,y\in X}Q(x,y)\pi(x)(f(y)-f(x))(g(y)-g(x)),$
${\mathrm{Ent}}(\rho)$ $:=\langle\rho,\log\rho\rangle_{\pi}=\sum_{x\in X}\rho(x)\log\rho(x)\pi(x).$
We say that $(X,Q,\pi)$ satisfies (MLSI) with optimal constant $\alpha_{MLSI}>0$ where
$\alpha_{MLSI}:=\inf_{\rho\in\mathcal{P}_{*}(X)}\frac{\mathcal{E}(\rho,\log\rho)}{\mathrm{Ent}(\rho)}.$
The following results hold for Bernoulli-Laplace models:
  1. (i)
    It was shown in [CDPP-09, Theorem 5.2] that (MLSI) holds with
    $\alpha_{MLSI}\geq\min\lambda-(\max\lambda-\min\lambda).$
  2. (ii)
    Assuming $\max\lambda\leq 3\min\lambda$, it was shown in [FM-16, Thm. 4.7] that the entropic curvature is lower bounded by
    $K=\frac{\min\lambda}{2}-\frac{7}{8}\left(\max\lambda-\min\lambda\right).$
  3. (iii)
    Employing the general relation between entropic curvature and the MLSI constant for any finite Markov chain in [EM-12, Theorem 1.5], (ii) implies
    $\alpha_{MLSI}\geq 2K=\min\lambda-\frac{7}{4}(\max\lambda-\min\lambda),$
    which is slightly weaker than (i).
  4. (iv)
    It was discussed at the end of [CMS-25] that employing Ollivier Ricci curvature and sectional curvature gives better curvature (see Theorem 5.6 below) and MLSI constant estimates, which do not depend on $\max\lambda$. Namely, non-negative sectional curvature and (53) below imply the improved inequality
    $\alpha_{MLSI}\geq\inf_{x\sim y}K_{ORC}(x,y)\geq\frac{1}{L}\left((L-N-1)\lambda_{1}+\sum_{i=1}^{N+1}\lambda_{i}\right)\geq\min\lambda$
    by [munch2023ollivier, Theorem 4.4]. Here we assume the following ordering of the intensities
    $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{L}.$
    Note that even in the case $\min\lambda=0$, this lower bound is still meaningful, for example, in the case $L=2N$ and $\lambda_{i}=0$ for $i\leq L/4$ and $\lambda_{i}\geq 1$ for $i>L/4$, we have
    $\alpha_{MLSI}\geq\frac{1}{4}.$
Theorem 5.6.
Let $(X,Q,\pi)$ be a Laplace-Bernoulli model with $N$ particles and $L$ sites and corresponding intensities
$0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{L}.$
Then $(X,Q,\pi)$ has non-negative Ollivier sectional curvature $\inf_{x\sim y}K_{SEC}(x,y)\geq 0$ and we have
$\inf_{x\sim y}K_{ORC}(x,y)\geq\frac{1}{L}\left((L-N-1)\lambda_{1}+\sum_{i=1}^{N+1}\lambda_{i}\right).$ (53)
Proof.
We start with the computation of $K_{ORC}(x,y)$ where the labels on vertices $x$ and $y$ are $x=(\underbrace{1,1,...,1}_{N},\underbrace{0,0,...,0}_{L-N})$ and $y=(0,\underbrace{1,1,...,1}_{N},\underbrace{0,0,...,0}_{L-N-1})$. There are two types of vertices $z$ which are common neighbors of $x$ and $y$ (that is, $z\in S_{1}(x)\cap S_{1}(y)$):
  • $z=(0,\underbrace{1,1,...,1}_{N-1},0,0,0,...,0,1,0...,0)$ where the last $1$ is at position $N+2\leq i\leq L$. In this case, we have $Q(x,z)=\lambda_{1}$ and $Q(y,z)=\lambda_{N+1}$.
  • $z=(\underbrace{1,1,...,1,0,1,...,1}_{N+1},0,0,...,0)$ where the first 00 is at position $2\leq j\leq N$. In this case, we have $Q(x,z)=Q(y,z)=\lambda_{j}$.
For vertices $z$ which are not common neighbors of $x$ and $y$, there is a perfect matching $\Phi$ from $S_{1}(x)\setminus S_{1}(y)$ to $S_{1}(y)\setminus S_{1}(x)$, namely
$\Phi(\tilde{x}):=\tilde{x}+y-x,$
by applying the tuple-addition. Moreover, we have
$Q(x,\tilde{x})=Q(y,\Phi(\tilde{x})).$
Hence, there exists a transport plan $\gamma$ from $Q(x,\cdot)$ to $Q(y,\cdot)$ with maximum transport distance $1$, which implies non-negative sectional curvature $K_{SEC}(x,y)\geq 0$. This transport plan is optimal and given by
$\gamma(\tilde{x},\tilde{y})=\begin{cases}Q(x,y),&\text{if $\tilde{x}=\tilde{y}=y$,}\\ Q(y,x),&\text{if $\tilde{x}=\tilde{y}=x$,}\\ Q(x,\tilde{x})\wedge Q(y,\tilde{y}),&\text{if $\tilde{x}=\tilde{y}\in S_{1}(x)\cap S_{1}(y)$,}\\ Q(x,\tilde{x}),&\text{if $\tilde{x}\in S_{1}(x)\setminus S_{1}(y)$ and $\tilde{y}=\Phi(\tilde{x})$,}\\ Q(x,\tilde{x})-(Q(x,\tilde{x})\wedge Q(y,\tilde{y}))&\text{if $\tilde{x}\in S_{1}(x)\cap S_{1}(y)$ and $\tilde{y}=y$,}\\ Q(y,\tilde{y})-(Q(x,\tilde{x})\wedge Q(y,\tilde{y}))&\text{if $\tilde{x}=x$ and $\tilde{y}\in S_{1}(x)\cap S_{1}(y)$,}\\ c,&\text{if $\tilde{x}=x$ and $\tilde{y}=y$,}\\ 0,&\text{otherwise.}\end{cases}$
Here, $c\geq 0$ is chosen such that $\sum_{\tilde{x},\tilde{y}\in X}\gamma(\tilde{x},\tilde{y})=1$. The Ollivier-Ricci curvature is therefore the total mass transported by distance 00, that is,
$K_{ORC}(x,y)$ $=Q(x,y)+Q(y,x)+\sum_{z\in S_{1}(x)\cap S_{1}(y)}Q(x,z)\wedge Q(y,z)$
$=\frac{1}{L}\left(\lambda_{1}+\lambda_{N+1}+(L-N-1)\min\lambda+\sum_{j=2}^{N}\lambda_{j}\right)$
$=\frac{1}{L}\left((L-N-1)\lambda_{1}+\sum_{i=1}^{N+1}\lambda_{i}\right).$
Analogously, we can compute Ollivier-Ricci curvature of any edge $u\sim v$ as
$K_{ORC}(u,v)$ $=\frac{1}{L}\left((L-N-1)\lambda_{\sigma(1)}+\sum_{i=1}^{N+1}\lambda_{\sigma(i)}\right)$
$\geq\frac{1}{L}\left((L-N-1)\lambda_{1}+\sum_{i=1}^{N+1}\lambda_{i}\right),$
where $\sigma\in\mathcal{S}_{L}$ is a suitable permutation. ∎
@@ 6 Some open questions
  • Do all Ricci flat or (S)-Ricci flat graphs, as defined in Definition 3.9, have non-negative entropic curvature, as it is the case for Ollivier Ricci and Bakry-Émery curvature?
@@
  • For the entropic curvature estimate of the Bernoulli-Laplace model, the article [FM-16] required the restriction $\frac{\max\lambda}{\min\lambda}\leq 3$. It would be interesting to investigate both entropic curvature and Bakry-Émery curvature in the case when $\frac{\max\lambda}{\min\lambda}$ is large.
@@
[…]
Yau Mathematical Sciences Center, Tsinghua University, Beijing, China, and
Department of Mathematics and Computer Science, Faculty of Science
Chulalongkorn University, Bangkok, Thailand
E-mail address: supanat.k@chula.ac.th
School of Mathematical Sciences, University of Science and Technology of China
Hefei, China
E-mail address: spliu@ustc.edu.cn
MPI MiS Leipzig, Leipzig, Germany, and
Faculty of Mathematics and Computer Science, Leipzig University, Leipzig, Germany
E-mail address: cfmuench@gmail.com
Department of Mathematical Sciences, Durham University, Durham, United Kingdom
E-mail address: norbert.peyerimhoff@durham.ac.uk